please help me with this equations:1/2(x+8)=7

Answer by Died2Soon(6)

First you need to remove the fraction to make the problem easier.
to remove the fraction multiply both sides by the denominator: {1/2(x+8)}2=7*2
you should get: x+8=14 because the 1/2 gets cancelled out by multiplying by 2
now subtract 8 from both sides: x+8-8=14-8
your answer should be: x=6
Hope this helps
=)